OpenStudy (anonymous):

Solve each equation on the interval 0

OpenStudy (anonymous):

\[\sin (3\Theta) = - 1\]

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

\[0 \le \theta < 2\pi\]

OpenStudy (anonymous):

@TheSmartOne you there? Don't want to annoy you or anything but I could use some help

OpenStudy (chaise):

Take the inverse sin of both sides, and then divide by 3 to solve this equation, remember that there will be four solutions to this equation between 0 and 2 pi. The easiest way to understand this is to draw the sides on a cartesian plane, recalling that: sin(x)=opp/hyp

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

when does sin(u) = -1?

OpenStudy (amistre64):

this is one of those basic trig values that youll have to memorize

OpenStudy (anonymous):

-90

OpenStudy (amistre64):

yes, or in the interval defined; 3pi/2

OpenStudy (amistre64):

so for what x does: 3x = 3pi/2 ?

OpenStudy (anonymous):

pi/2